OpenStudy (anonymous):

Evaluate the Integral\[\int\limits_{C}^{}ydx+xdy\] C is the curve\[y=x^3, 0 \le x \le 3\]

OpenStudy (anonymous):

Parameterize \(C\) with \[\begin{cases} x=t\\ y=t^3\\ 0\le t\le3 \end{cases}~~\implies~~\begin{cases}dx=dt\\dy=3t^2~dt\end{cases}\] Then \[\int_C y~dx+x~dy=\int_0^3\left(t^3+t(3t^2)\right)~dt\]

OpenStudy (anonymous):

I must've done something weird. I thought I did that but I was getting a weird answer. Now I seem to be getting the correct one. Thank You.